# The quark gap equation in light-cone gauge

### v1

@@ Abstract
We calculate the quark self-energy correction in light-cone gauge motivated by distribution amplitudes whose definition implies a Wilson line. The latter serves to preserve the gauge invariance of the hadronic amplitudes and becomes trivial in light-cone gauge. Therefore, the calculation of the distribution amplitudes simplifies significantly provided that wave functions and propagators are obtained in that gauge. In here, we explore the corresponding Dyson-Schwinger equation in its leading truncation and with a dressed vertex derived from a Ward identity in light-cone gauge. The quark’s mass and wave renormalization functions, as well as a third complex-valued amplitude, are found to depend on the relative orientation of the quark momentum and a light-like four-vector, which expresses a geometric gauge dependence of the propagator.
@@ I Introduction
Noncovariant gauges have a long history and, as their name implies, one of their features is the breaking of relativistic covariance. A typical example is provided by the light-cone gauge, which is a noncovariant physical gauge defined by,
@@
where $A^{a}_{\mu}(x)$ is a massless Yang-Mills field and $n_{\mu}$ is a light-like four-vector defined by $n^{2}=0$. Since the four-vector $n$ defines a preferential axis in space-time, the above condition is more generally referred to as axial gauge. Despite the lack of covariance, the strong interest in axial gauges can be attributed to the decoupling of the Faddeev-Popov ghosts from the gauge field, thereby eliminating the unphysical degrees of freedom in the theory Leibbrandt:1987qv . This is because the ghost-gluon vertex is proportional to the vector $n$ which projects out the gluon fields due to the gauge condition in Eq. (1). Consequently, $n^{\mu}D_{\mu\nu}^{ab}=0$ in any Feynman diagram, where $D_{\mu\nu}^{ab}$ is the gluon propagator in light-cone gauge. This observation alone may be of interest to the nonperturbative calculation of the gluon propagator Aguilar:2006gr ; Aguilar:2008xm , as the gluon also decouples from the ghost in the Dyson-Schwinger equation (DSE).
@@
×$\times\,\langle M|\bar{\psi}_{q}(0)\gamma\cdot n\,W(0,\lambda n)\psi_{q}(\lambda n)|M\rangle\,,$ (2)
@@
[…]
in which $\mathcal{P}$ denotes the path-ordering operator. In light-cone gauge (1) this operator is trivial, which is why the calculation of distributions is attractive in this gauge. However, this comes at a price: the necessity to calculate nonperturbative quark propagators in light-cone gauge. In a functional continuum approach to Quantum Chromodynamics (QCD), this is commonly done by solving the respective DSEs of the fermion and gauge fields in Landau or covariant $R_{\xi}$ gauges Lessa:2022wqc .
@@ II Gluon propagator in light-cone gauge
$\mathcal{L}_{\text{fix}}=-\frac{1}{2\alpha}\left(n\cdot A^{a}\right)^{2},$ (4)
@@
[…]
Therefore, the location of the $(n\cdot q)^{-1}$ pole is not unique. This ambiguity can be addressed by choosing $n=(|\vec{n}|,\vec{n})$ and introducing a dual vector $n^{*}=(|\vec{n}|,-\vec{n})$, $n^{*2}=0$, $n^{*}\cdot n>0$. According to Ref. Leibbrandt:1983pj , the ML prescription is then given by,
@@
$\Sigma(p,n,n^{*})$ $=a\,i\gamma\cdot p+b\,\mathbb{1}+c\,i\gamma\cdot n$
@@
$\gamma\cdot n\,\gamma$ $\cdot n^{*}\gamma\cdot p+\gamma\cdot p\,\gamma\cdot n^{*}\gamma\cdot n$
$=$ $\ (2\,n^{*}\!\cdot p)\gamma\cdot n-(2\,n\cdot p)\gamma\cdot n^{*}\!+(2\,n^{*}\!\!\cdot n)\gamma\cdot p\,,$ (13)
which simplifies Eq. (12) to:
$\Sigma(p,n,n^{*})=A\,i\gamma\cdot p+B\,\mathbb{1}+C\,i\gamma\cdot n+D\,i\gamma\cdot n^{*}\,.\$ (14)
The structure of Eq. (14) can be understood within the framework of the Newman-Penrose tetrad scheme Newman:1961qr , in which any four-dimensional vector can be represented in terms of four null vectors, and is discussed in detail in Ref. Leibbrandt:1984be .
Appropriate projections allow to obtain integral equations for all four amplitudes, $A,B,C$ and $D$, and in a one-loop calculation they can be solved with the usual Feynman parametrizations, though the four equations remain uncoupled. In a nonperturbative functional approach based on the DSE, the amplitudes are described by a set of nonlinear coupled integral equations. We numerically solve this system of equations in light-cone gauge and discuss their solutions in the following sections.
@@ III Gap equation in light-cone gauge
For the phenomenological reasons expounded in Section I, we are interested in nonperturbative quark propagators which are solutions of the quark gap equation. To this end, we solve the equation of motion of the quark, which in quantum field theory is known as DSE. For the quark of a given flavor the inverse quark propagator is obtained from111 We employ Euclidean metric, $\left\{\gamma_{\mu},\gamma_{\nu}\right\}=2\delta_{\mu\nu}$, with hermitian Dirac matrices: $\gamma^{\dagger}_{\mu}=\gamma_{\mu}$. Furthermore, $\gamma_{5}=\gamma_{5}^{\dagger}=\gamma_{4}\gamma_{1}\gamma_{2}\gamma_{3}$, with $\mathrm{Tr}\left[\gamma_{5}\gamma_{\mu}\gamma_{\nu}\gamma_{\alpha}\gamma_{\beta}\right]=-4\epsilon_{\mu\nu\alpha\beta}$, $\sigma_{\mu\,\nu}=\frac{i}{2}\left[\gamma_{\mu},\gamma_{\nu}\right]$ and a space-like vector $p_{\mu}$ is characterized by $p^{2}>0$.,
@@
where $q=k-p$ is the gluon momentum, $Z_{1}(\mu,\Lambda)$, $Z_{2}(\mu,\Lambda)$ and $Z_{4}(\mu,\Lambda)$ are the vertex, wave function and mass renormalization constants, respectively, $m(\mu)$ is the renormalized current-quark mass and $t^{a}=\lambda^{a}/2$ are the SU(3) group generators. In the self-energy integral, $\Lambda$ is a Poincaré-invariant cut-off while $\mu$ is the renormalization point chosen such that $\Lambda\gg\mu$. For a review of phenomenological applications of the DSE, we refer to Ref. Bashir:2012fs .
The quark propagator is a gauge-dependent Green function and if we specify the light-front gauge, the gluon propagator must be of the form in Eq. (5). The gluon obeys its own DSE, but no known solution exists in light-cone gauge. We assume a nonperturbative gluon propagator,
$D_{\mu\nu}^{ab}(q)=\delta^{ab}\,\frac{\Delta(q^{2})}{q^{2}}\left(\delta_{\mu\nu}-\frac{n_{\mu}q_{\nu}+n_{\nu}q_{\mu}}{n\cdot q}\right),$ (16)
with a dressing function $\Delta(q^{2})$ for which we will define a model. Taking into account Eq. (14), the solutions of the DSE (15) are generally written as,
@@
The propagator differs from its form in covariant gauges not merely because of two additional scalar amplitudes, but also due the dependence of all scalar amplitudes on the relative orientation of the four-vectors $p$, $n$ and $n^{*}$. We will see that this is a signature of the light-cone gauge dependence.
@@
where the form factors $\lambda_{i}\equiv\lambda_{i}(k,p,n,n^{*})$ are known to one-loop Leibbrandt:1983zd . However, since the form of Eq. (18) is derived in perturbation theory, the corresponding Ward identity is valid for propagators with constant quark mass, but not for the nonperturbative solutions in Eq. (17).
We therefore generalize the vertex in Eq. (18) and in analogy with the standard Ball-Chiu vertex we add the two components, $\gamma\cdot k\,\gamma\cdot p\,(k+p)_{\mu}$ and $(k+p)_{\mu}$. Thus, our ansatz is given by the non-transverse vertex decomposition,
$\Gamma_{\mu}(k,p,n,n^{*})$ $=\lambda_{1}\gamma_{\mu}+\lambda_{2}\left(\gamma\cdot k+\gamma\cdot p\right)(k+p)_{\mu}$
@@
$\lambda_{1}(k,p,n,n^{*})$ $=\frac{A(k)+A(p)}{2}\ ,$ (23)
$\lambda_{2}(k,p,n,n^{*})$ $=\frac{1}{2}\frac{A(k)-A(p)}{k^{2}-p^{2}}\ ,$ (24)
@@
where a dependence of the scalar amplitudes on $n$ and $n^{*}$ is implicit.
If the gluon propagator and quark-gluon vertex were bare, the arguments presented in Section II that justify a Wick rotation would be valid provided the quark-mass function was constant, with poles located in the second and fourth quadrants of the complex plane. This, however, is not the case as the nonperturbative mass function, defined by $M(p^{2})=B(p^{2})/A(p^{2})$, is characterized by complex-conjugate poles or branch cuts El-Bennich:2016qmb . An interpolation formalism from instant-frame to light-front dynamics in a QCD model in $1+1$ dimensions and infinite number of colors from was shown to allow for a matching of Minkowski and Euclidean space Ma:2021yqx , though the case of QCD in $3+1$ dimensions beyond perturbative studies is an open question. A detailed study of the singularity structure of propagators in the Bethe-Salpeter equation demonstrated that the light-front wave function calculated with light-front coordinates defined in Euclidean metric is identical to the one in Minkowski metric for the case of a monopole model of the Bethe-Salpeter amplitude Eichmann:2021vnj . Eschewing these conceptual difficulties, we here make bold to work directly in Euclidean space.
@@
With these projections we arrive at a set of coupled integral equations which in rainbow truncation, $\Gamma_{\mu}=Z_{2}\gamma_{\mu}$ Bloch:2002eq , is given by,
$A(p,n,n^{*})$ $=\,Z_{2}+\frac{2D(p,n,n^{*})}{n\cdot p}-\frac{iZ_{2}^{2}g^{2}C_{f}}{4\,n\cdot p}\!\int^{\Lambda}\!\frac{d^{4}k}{(2\pi)^{4}}\,D_{\mu\nu}(q)\operatorname{Tr}_{\mathrm{CD}}\left[\gamma\cdot n\,\gamma_{\mu}S(k,n,n^{*})\gamma_{\nu}\right],$ (32)
$B(p,n,n^{*})$ $=\,Z_{4}m(\mu)+\frac{Z_{2}^{2}g^{2}C_{f}}{4}\!\int^{\Lambda}\!\frac{d^{4}k}{(2\pi)^{4}}\,D_{\mu\nu}(q)\operatorname{Tr}_{\mathrm{CD}}\left[\gamma_{\mu}S(k,n,n^{*})\gamma_{\nu}\right],$ (33)
$C(p,n,n^{*})$ $=\frac{p^{2}}{n\cdot p}[Z_{2}-A(p,n,n^{*})]-D(p,n,n^{*})\frac{n^{*}\!\cdot p}{n\cdot p}-\frac{iZ_{2}^{2}g^{2}C_{f}}{4\,n\cdot p}\!\int^{\Lambda}\!\!\!\frac{d^{4}k}{(2\pi)^{4}}\,D_{\mu\nu}(q)\operatorname{Tr}_{\mathrm{CD}}[\gamma\cdot p\,\gamma_{\mu}S(k,n,n^{*})\gamma_{\nu}]\,,$ (34)
$D(p,n,n^{*})$ $=\frac{p^{2}}{n^{*}\!\cdot p}[Z_{2}-A(p,n,n^{*})]-C(p,n,n^{*})\frac{n\cdot p}{n^{*}\!\cdot p}-\frac{iZ_{2}^{2}g^{2}C_{f}}{4\,n^{*}\!\cdot p}\!\!\int^{\Lambda}\!\!\!\frac{d^{4}k}{(2\pi)^{4}}\,D_{\mu\nu}(q)\operatorname{Tr}_{\mathrm{CD}}[\gamma\cdot p\,\gamma_{\mu}S(k,n,n^{*})\gamma_{\nu}]\,.$ (35)
@@
Closer inspection of Eqs. (34) and (35), multiplying them by $n\cdot p$ and $n^{*}\!\cdot p$ respectively, reveals that they are identical. Indeed, one of the functions, $C(p,n,n^{*})$ or $D(p,n,n^{*})$, is superfluous and we shall see that $D(p,n,n^{*})\equiv 0$. This is not surprising, as the introduction of $n^{*}$ is an artifact to tame divergences that stem from the the $(n\cdot p)^{-1}$ factor in the gluon propagator (16). In Euclidean space, though, the ML prescription is not necessary in solving the DSE (15).
We make the replacement,
$Z_{2}^{2}\,g^{2}\Delta(q^{2})/q^{2}=Z_{2}^{2}\,\alpha(q^{2})/q^{2}\,,$ (36)
which introduces a model interaction, $\alpha(q^{2})$, for the gluon dressing Maris:1999nt ,
$\frac{\alpha(q^{2})}{q^{2}}=\frac{4\pi^{2}}{\omega^{6}}Dq^{2}e^{-q^{2}\!/\omega^{2}}\!+\frac{8\pi^{2}\gamma_{m}\mathcal{G}(q^{2})}{\ln\!\Big{[}\tau+\big{(}1+q^{2}/\Lambda_{\mathrm{QCD}}^{2}\big{)}^{\!2}\Big{]}},$ (37)
in which $\gamma_{m}=12/(33-2N_{f})$, $N_{f}=4$, $\Lambda_{\mathrm{QCD}}=0.234$ GeV, $\tau=e^{2}-1$ and $\mathcal{G}(q^{2})=\left[1-\exp(-q^{2}/4m_{t}^{2})\right]/q^{2}$ with $m_{t}=0.5$ GeV. Computing the traces yields the three integral equations,
$A(p,n)=$ $\ Z_{2}+\frac{2Z_{2}^{2}C_{f}}{n\cdot p}\int^{\Lambda}\!\!\frac{d^{4}k}{(2\pi)^{4}}\frac{\alpha(q^{2})}{q^{2}}\,\sigma_{A}(k,n)\,n\cdot k\,,$ (38)
@@
$+$ $\ 2Z_{2}^{2}C_{f}\int^{\Lambda}\!\frac{d^{4}k}{(2\pi)^{4}}\frac{\alpha(q^{2})}{q^{2}}\,\sigma_{B}(k,n)\,,$ (39)
$C(p,n)=$ $\ \frac{p^{2}}{n\cdot p}\big{[}Z_{2}-A(p,n)\big{]}$
$+$ $\ 2Z_{2}^{2}C_{f}\int^{\Lambda}\!\frac{d^{4}k}{(2\pi)^{4}}\frac{\alpha(q^{2})}{q^{2}}\Big{[}\sigma_{C}(k,n)$
@@
with the scalar dressing functions,
@@
Figure 1: The quark-dressing functions $A(p,n)$ and $B(p,n)$ calculated with a bare vertex and with the WFGTI ansatz of Eq. (19).
@@
$B(\mu,n)=m(\mu,n)=5\;\mathrm{MeV}\quad\text{and}\quad A(\mu,n)=1\ ,$ (45)
at a renormalization scale $\mu=19$ GeV, whereas $C(k,n)$ is left unconstrained. [^2] We can also impose the renormalization condition $C(\mu=0,n)=0$, as suggested by a one-loop calculation in Ref. Mirja2020 , though the unconstrained numerical solution tends to small $C(p,n)$ values in the limit $p^{2}\rightarrow 0$.
Likewise, we derive the integral equations (28) to (31) for the vertex in Eq. (19), the expressions of which we here omit due to their length.
Figure 3: Real and imaginary parts of the dressing function $C(p,n)$ obtained with the RL and WFGTI vertices.

### v2

@@ Abstract
An exploratory nonperturbative calculation of the quark propagator in light-cone gauge is motivated by distribution amplitudes whose definition implies a Wilson line. The latter serves to preserve the gauge invariance of the hadronic amplitudes and becomes trivial in light-cone gauge. To that end, we explore the corresponding Dyson-Schwinger equation in its leading truncation and with a dressed vertex derived from a Ward identity in light-cone gauge. The quark’s mass and wave renormalization functions are found to depend on the orientation of the quark momentum relative to the light-like four-vector below 1 GeV, which expresses the light-cone gauge dependence of the propagator, while a third, complex-valued amplitude exhibits little dependence on that orientation and vanishes in the quark’s rest frame.
@@ I Introduction
Noncovariant gauges have a long history and, as their name implies, one of their features is the breaking of relativistic covariance. A typical example is provided by the light-cone gauge, which is a physical gauge defined by,
@@
where $A^{a}_{\mu}(x)$ is a massless Yang-Mills field and the light-like four-vector defined by $n^{2}=0$. Since the four-vector $n$ defines a preferential axis in space-time, the above condition is more generally referred to as axial gauge. Despite the lack of covariance, the strong interest in axial gauges can be attributed to the decoupling of the Faddeev-Popov ghosts from the gauge field, thereby eliminating the unphysical degrees of freedom in the theory Leibbrandt:1987qv . This is because the ghost-gluon vertex is proportional to the vector $n$ which projects out the gluon fields due to the gauge condition in Eq. (1). Consequently, $n^{\mu}D_{\mu\nu}^{ab}=0$ in any Feynman diagram, where $D_{\mu\nu}^{ab}$ is the gluon propagator in light-cone gauge. This observation alone may be of interest to the nonperturbative calculation of the gluon propagator Aguilar:2006gr ; Aguilar:2008xm , as the gluon also decouples from the ghost in the Dyson-Schwinger equation (DSE).
@@
$\times\,\langle M|\bar{\psi}_{q}(0)\gamma\cdot n\,W(0,\lambda n)\psi_{q}(\lambda n)|M\rangle\,$, (2)
@@
$W(0,n\lambda)=\mathcal{P}e^{-ig\int_{\lambda}^{0}n\,\cdot A(n\xi)d\xi}\,$, (3)
in which $\mathcal{P}$ denotes the path-ordering operator. In light-cone gauge (1) this operator is trivial, which is why the calculation of distributions is attractive in this gauge. However, this comes at a price: the necessity to compute nonperturbative quark propagators in light-cone gauge. In a functional continuum approach to Quantum Chromodynamics (QCD), however, this is commonly done by solving the respective DSEs of the fermion and gauge fields in Landau or covariant $R_{\xi}$ gauges Lessa:2022wqc .
@@ II Gluon propagator in light-cone gauge
$\mathcal{L}_{\text{fix}}=-\frac{1}{2\alpha}\left(n\cdot A^{a}\right)^{2}.$ (4)
@@
$n_{0}^{2}-\vec{n}^{\,2}=0\ \ \Longrightarrow\ \ n_{0}=\pm|\vec{n}|\,.$ (7)
Therefore, the location of the $(n\cdot q)^{-1}$ pole is not unique. This ambiguity can be addressed by choosing $n=(|\vec{n}|,\vec{n})$ and introducing a dual vector $n^{*}=(|\vec{n}|,-\vec{n})$, $^{*2}=0, nn$$^{*}\cdot n>0$. According to Ref. Leibbrandt:1983pj , the ML prescription is then given by,
@@
$\hskip-5.69054pt\Sigma(p,n,n^{*})$ $=a\,i\gamma\cdot p+b\,\mathbb{1}+c\,i\gamma\cdot n$
@@
$\hskip-2.84526pt\gamma\cdot n\,\gamma$ $\cdot n^{*}\gamma\cdot p+\gamma\cdot p\,\gamma\cdot n^{*}\gamma\cdot n$
$=$ $\ (2\,n^{*}\!\cdot p)\gamma\cdot n-(2\,n\cdot p)\gamma\cdot n^{*}\!+(2\,n^{*}\!\!\cdot n)\gamma\cdot p\,.$ (13)
This simplifies Eq. (12) which becomes:
$\Sigma(p,n,n^{*})=A\,i\gamma\cdot p+B\,\mathbb{1}+C\,i\gamma\cdot n+D\,i\gamma\cdot n^{*}\,$,\\ (14)
where all four dressing functions are functions of $p$, $n$ and $n^{*}$. The structure of Eq. (14) can be understood within the framework of the Newman-Penrose tetrad scheme Newman:1961qr , in which any four-dimensional vector can be represented as a superposition of four null-vectors, and is discussed in detail in Ref. Leibbrandt:1984be .
Appropriate projections allow to obtain integral equations for all four amplitudes, $A,B,C$ and $D$, and in a one-loop calculation they can be solved with the usual Feynman parametrizations, though the four equations remain uncoupled. On the other hand, in a nonperturbative functional approach solving the DSE of the quark, the amplitudes are described by a set of nonlinear coupled integral equations. We solve this system of equations in light-cone gauge and discuss their numerical solutions in the following sections.
@@ III Gap equation in light-cone gauge
For the phenomenological reasons expounded in Section I, we are interested in nonperturbative quark propagators which are solutions of the quark gap equation. To this end, we solve the equation of motion of the quark, which in quantum field theory is known as DSE. For the quark of a given flavor the inverse quark propagator is obtained from [^1] We employ Euclidean metric, $\left\{\gamma_{\mu},\gamma_{\nu}\right\}=2\delta_{\mu\nu}$, with hermitian Dirac matrices: $\gamma^{\dagger}_{\mu}=\gamma_{\mu}$. Furthermore, $\gamma_{5}=\gamma_{5}^{\dagger}=\gamma_{4}\gamma_{1}\gamma_{2}\gamma_{3}$, with $\mathrm{Tr}\left[\gamma_{5}\gamma_{\mu}\gamma_{\nu}\gamma_{\alpha}\gamma_{\beta}\right]=-4\epsilon_{\mu\nu\alpha\beta}$, $\sigma_{\mu\,\nu}=\frac{i}{2}\left[\gamma_{\mu},\gamma_{\nu}\right]$ and a space-like vector $p_{\mu}$ is characterized by $p^{2}>0$.,
@@
where q=kpq=k-p is the gluon momentum, $Z_{1}(\mu,\Lambda)$, $Z_{2}(\mu,\Lambda)$ and $Z_{4}(\mu,\Lambda)$ are the vertex, wave function and mass renormalization constants, respectively, $m(\mu)$ is the renormalized current-quark mass and $t^{a}=\lambda^{a}/2$ are the SU(3) group generators. In the self-energy integral, $\Lambda$ is a Poincaré-invariant cut-off while $\mu$ is the renormalization point chosen such that $\Lambda\gg\mu$. For a review of phenomenological applications of the DSE, we refer to Ref. Bashir:2012fs .
The quark propagator is a gauge-dependent Green function and if we specify the light-front gauge, the gluon propagator must be of the form in Eq. (5). Generalizing, the nonperturbative gluon propagator can be written as Cornwall:1981zr ,
$D_{\mu\nu}^{ab}(q)=\delta^{ab}\,\Delta(q^{2})\left(\delta_{\mu\nu}-\frac{n_{\mu}q_{\nu}+n_{\nu}q_{\mu}}{n\cdot q}\right),$ (16)
where $\Delta(q^{2})$ is the gluon’s dressing function which we introduce in Eq. (36). Taking into account Eq. (14), the solutions of the DSE (15) are generally written as,
@@
The propagator differs from its form in covariant gauges not merely because of two additional scalar amplitudes, but also due the dependence of all scalar functions on the relative orientation of the four-vectors $p$, $n$ and $n^{*}$. We will see that this is a signature of the light-cone gauge dependence.
@@
where the form factors $\lambda_{i}\equiv\lambda_{i}(k,p,n,n^{*})$ are known to one-loop Leibbrandt:1983zd . However, since the form of Eq. (18) is derived in perturbation theory, the corresponding Ward identity is valid for propagators with constant quark mass, but not for the nonperturbative solutions of Eq. (17).
We therefore generalize the vertex in Eq. (18) and in analogy with the standard Ball-Chiu vertex we add the two components, $\gamma\cdot(k+p)(k+p)$ and $(k+p)$. Thus, combining the two terms proportional to $n_{\mu}$ in Eq. (18), our ansatz is given by the non-transverse vertex decomposition,
$\Gamma_{\mu}(k,p,n,n^{*})$ $=\lambda_{1}\gamma_{\mu}+\lambda_{2}\,\gamma\cdot\left(k+p\right)(k+p)_{\mu}$
@@
$\lambda_{1}(k,p,n,n^{*})$ $=\frac{A(k))+A(p,))}{2}\ ,$ (23)
$\lambda_{2}(k,p,n,n^{*})$ $=\frac{1}{2}\frac{A(k))-A(p)}{k^{2}-p^{2}}\ ,$ (24)
@@
where a dependence of the scalar amplitudes $A,B,C$ and $D$ on the light-like vector $n$ and $n^{*}$ is implicit.
If the gluon propagator and quark-gluon vertex were bare, the arguments presented in Section II that justify a Wick rotation would be valid, provided a constant quark-mass function with poles located in the second and fourth quadrants of the complex plane. This, however, is not the case as the nonperturbative mass function, defined by $M(p^{2})=B(p^{2})/A(p^{2})$, is characterized by complex-conjugate poles or branch cuts El-Bennich:2016qmb . An interpolation formalism from instant-frame to light-front dynamics in a QCD model in $1+1$ dimensions and infinite number of colors was shown to allow for a matching of Minkowski and Euclidean space Ma:2021yqx , though the case of QCD in $3+1$ dimensions beyond perturbative studies is an open question. A detailed study of the singularity structure of propagators in the Bethe-Salpeter equation demonstrated that the light-front wave function calculated with light-front coordinates defined in Euclidean metric is identical to the one in Minkowski metric for the case of a monopole model of the Bethe-Salpeter amplitude Eichmann:2021vnj , while the DSE in Minkowski space has been treated, for instance, in Refs. Sauli:2002tk ; Duarte:2022yur . Eschewing these conceptual difficulties, we here make bold to work directly in Euclidean space.
@@
With these projections we arrive at a set of coupled integral equations which is given in rainbow truncation by,
$A(p,n,n^{*})$ $=\,Z_{2}+\frac{2D(p,n,n^{*})}{n\cdot p}-\frac{iZ_{2}g^{2}C_{f}}{4\,n\cdot p}\!\int^{\Lambda}\!\frac{d^{4}k}{(2\pi)^{4}}\,D_{\mu\nu}(q)\operatorname{Tr}_{D}\left[\gamma\cdot n\,\gamma_{\mu}S(k,n,n^{*})\gamma_{\nu}\right],$ (32)
$B(p,n,n^{*})$ $=\,Z_{4}m(\mu)+\frac{Z_{2}g^{2}C_{f}}{4}\!\int^{\Lambda}\!\frac{d^{4}k}{(2\pi)^{4}}\,D_{\mu\nu}(q)\operatorname{Tr}_{D}\left[\gamma_{\mu}S(k,n,n^{*})\gamma_{\nu}\right],$ (33)
$C(p,n,n^{*})$ $=\frac{p^{2}}{n\cdot p}[Z_{2}-A(p,n,n^{*})]-D(p,n,n^{*})\frac{n^{*}\!\cdot p}{n\cdot p}-\frac{iZ_{2}g^{2}C_{f}}{4\,n\cdot p}\!\int^{\Lambda}\!\!\!\frac{d^{4}k}{(2\pi)^{4}}\,D_{\mu\nu}(q)\operatorname{Tr}_{D}[\gamma\cdot p\,\gamma_{\mu}S(k,n,n^{*})\gamma_{\nu}]\,$, (34)
$D(p,n,n^{*})$ $=\frac{p^{2}}{n^{*}\!\cdot p}[Z_{2}-A(p,n,n^{*})]-C(p,n,n^{*})\frac{n\cdot p}{n^{*}\!\cdot p}-\frac{iZ_{2}g^{2}C_{f}}{4\,n^{*}\!\cdot p}\!\!\int^{\Lambda}\!\!\!\frac{d^{4}k}{(2\pi)^{4}}\,D_{\mu\nu}(q)\operatorname{Tr}_{D}[\gamma\cdot p\,\gamma_{\mu}S(k,n,n^{*})\gamma_{\nu}]\,.$ (35)
@@
Closer inspection of Eqs. (34) and (35), multiplying them by $n\cdot p$ and $n^{*}\!\cdot p$ respectively, reveals that they are identical. Indeed, one of the functions, either $C(p,n,n^{*})$ or $D(p,n,n^{*})$, is superfluous and it turns out that $D(p,n,n^{*})\equiv 0$. This is not surprising, as the introduction of $n^{*}$ is an artifact to tame divergences stemming from the $(n\cdot q)^{-1}$ factor in the gluon propagator (16). As we will see, the poles occurring in the denominators of the DSE (15) can be dealt with differently.
For the nonperturbative gluon propagator, that incorporates the correct ultraviolet behavior and is renormalizable, we employ the propagator derived by Cornwall within the framework of pinch technique Cornwall:1981zr ,
$\Delta^{-1}(q^{2})=\left[q^{2}+m_{g}^{2}(q^{2})\right]\!bg^{2}\ln\left[\frac{q^{2}+4m_{g}^{2}(q^{2})}{\omega^{2}}\right],$ (36)
with the running dynamical gluon mass,
$m_{g}^{2}(q^{2})=m_{g}^{2}\left(\frac{\ln\left[(q^{2}+4m_{g}^{2})/\omega^{2}\right]}{\ln\left[4m_{g}^{2}/\omega^{2}\right]}\right)^{\!\!-\tfrac{12}{11}}\!\!.$ (37)
The relationship between $q$, $\omega$ and $\mu$ is defined by $\Delta^{-1}(q^{2}=\mu^{2})=1/\mu^{2}$ aside of mass terms, and we remark that the product $g^{2}\Delta(q^{2})$ is independent of the strong coupling. The values for $b$, $m_{g}$ and $\omega$ will be discussed in Section IV. We remind Cornwall’s argument that the full, renormalized propagator should remain a function of Lorentz-invariant quantities only, and that in light-cone gauge a scalar Green function of one momentum cannot depend on $n_{\mu}$ Cornwall:1981zr ; Cornwall:1974hz ; thus, the dressing function $\Delta(q^{2})$ only depends on $q^{2}$, whereas the light-like vector appears in the free propagator (5).
We note that a contemporary understanding Binosi:2014aea of the pinch technique led to the development of a renormalization-group-invariant (RGI) running interaction, computed via a combination of DSE- and lattice-QCD results in Ref. Aguilar:2009nf . We employed both, the RGI interaction and the Cornwall propagator in Eq. (36), to solve the integral equations (38), (39) and (40), and found minimal qualitative and quantitative differences for the scalar functions $A(p,n)$, $B(p,n)$ and $C(p,n)$. The numerical results presented in Sec. IV were obtained with the pinch-technique gluon propagator of Eq. (36).
After taking the Dirac traces, we are left with three coupled integral equations,
$A(p,n)=$ $\ Z_{2}+\frac{2Z_{2}C_{f}}{n\cdot p}\!\int^{\Lambda}\!\!\frac{d^{4}k}{(2\pi)^{4}}g^{2}\Delta(q^{2})\,\sigma_{A}(k,n)\,n\cdot k\,$, (38)
@@
$+$ $\ 2Z_{2}C_{f}\int^{\Lambda}\!\frac{d^{4}k}{(2\pi)^{4}}g^{2}\Delta(q^{2})\,\sigma_{B}(k,n)\,$, (39)
$C(p,n)=$ $\ Z_{c}+\frac{p^{2}}{n\cdot p}\big[Z_{2}-A(p,n)\big]$
$+$ $\ 2Z_{2}C_{f}\int^{\Lambda}\!\frac{d^{4}k}{(2\pi)^{4}}g^{2}\Delta(q^{2})\Big[\sigma_{C}(k,n)$
@@
where we define the scalar dressing functions,
@@
Figure 1: The solution $A(p,n)$ and $B(p,n)$ of the quark DSE obtained with a bare vertex and with the WFGTI ansatz of Eq. (19).
@@
$B(\mu,n)=m(\mu)=25\;\mathrm{MeV}\quad\text{and}\quad A(\mu,n)=1\ ,$ (45)
at the renormalization scale $\mu=2$ GeV, and we introduce the renormalization constant $Z_{c}$ in Eq. (40) to the effect that $C(\mu,n)=0$ at small momenta, as suggested by a one-loop calculation in Ref. Mirja2020 . The cut-off can be formally taken to infinity, though numerically we employ $\Lambda=100$ GeV and our DSE solutions are stable for larger values of $\Lambda$. Likewise, we derive the integral equations (28) to (31) for the dressed vertex ansatz in Eq. (19), the expressions of which we here omit due to their length.
Figure 3: Real and imaginary parts of the dressing function $C(p,n)$ obtained with the bare and WFGTI vertices.